# A Black Hole Airy Tail

### v3

@@ Abstract
In Jackiw-Teitelboim (JT) gravity, which is dual to a random matrix ensemble, the annealed entropy differs from the quenched entropy at low temperatures and goes negative. However, computing the quenched entropy in JT gravity requires a replica limit that is poorly understood. To circumvent this, we define an intermediate quantity called the semi-quenched entropy, which has the positivity properties of the quenched entropy, while requiring a much simpler replica trick. We compute this in JT gravity in different regimes using i) a bulk calculation involving wormholes corresponding to the Airy limit of the dual matrix integral and ii) a boundary calculation involving one-eigenvalue instanton saddles proposed by Hernández-Cuenca, demonstrating consistency between these two calculations in their common regime of validity. We also clarify why similar one-eigenvalue instanton saddles cannot be used to compute the quenched entropy due to a breakdown of the saddle-point approximation for the one-eigenvalue instanton in the replica limit.
@@ I Introduction
In recent years, JT gravity has emerged as a tool that has significantly enhanced our understanding of quantum gravity. As a two-dimensional toy model of quantum gravity, it is computationally tractable while still exhibiting interesting non-perturbative effects. The action of this theory is given by [1]
@@
In [2], the Euclidean gravitational path integral of JT gravity was computed exactly to all perturbative orders in $e^{-S_{0}}$, in the $S_{0}\to\infty$ limit. The resulting asymptotic expansion agrees with the double-scaling limit of a random matrix theory (RMT):
@@
whose potential $V(H)$ is specified by the leading-order density of states:
@@
Figure 1: Comparing semi-quenched, annealed and quenched second Rényi entropies, zoomed in near the Airy edge $\beta=\alpha e^{2S_{0}/3}$ (with $\gamma=1$). Semi-quenched and annealed Rényi’s are obtained analytically, whereas the quenched Rényi is obtained numerically. Inset: We zoom in on the quenched and semi-quenched entropies at low temperatures and plot them on a log-log scale. Notice the different low temperature behavior: the quenched entropy has a $\alpha^{-3}$ power-law decay which can be obtained analytically (dotted blue line), whereas the semi-quenched entropy falls off exponentially. Nevertheless, as expected (and unlike the annealed entropy), our calculation shows that both entropies remain positive.
@@
In the absence of matter fields, the only observable in pure JT gravity is the energy spectrum. The obvious quantity of interest is thus the thermal entropy, obtained from the thermal partition function $Z\left(\beta\right)$. However, in an ensemble of theories where one only computes averaged quantities, one has to be careful about which entropy is being computed, since the average can be taken in different ways. Conventionally, the quantities discussed in the literature are the annealed entropy and the quenched entropy. The annealed entropy $S_{A}\left(\beta\right)\equiv\left(1-\beta\partial_{\beta}\right)\log\langle Z\left(\beta\right)\rangle$ differs from the quenched entropy, $S_{Q}\left(\beta\right)\equiv\left(1-\beta\partial_{\beta}\right)\langle\log Z\left(\beta\right)\rangle$, which is truly the average entropy of the ensemble. An important distinction between these entropies is that the quenched entropy is positive, while the annealed entropy can be negative.
@@
In gravitational theories, the annealed entropy is much easier to compute. When the entropy is self-averaging, $S_{A}=S_{Q}$. However, in JT gravity, $S_{A}$ fails to be self-averaging at sufficiently low temperatures. The consequent mismatch between the two entropies was analyzed in [3], where they found $S_{A}<0$ at $\beta=O(e^{2S_{0}/3})$ [5]. The negativity of the annealed entropy is not unique to the JT toy model; rather, it persists in higher-dimensional theories for all near-extremal black holes that are not supersymmetric at extremality [6, 7, 8, 9]. From the gravity point of view, this low temperature regime is thus among the most promising to probe the physics of only a small number of black hole microstates, specifically, the ones lying close to the groundstate. It is thus interesting to understand the difference between these entropies in this regime.
However, calculating the quenched entropy is not easy. To compute the quenched entropy, one uses the so-called no-replica trick: $\langle\log Z(\beta)\rangle=\partial_{m}\langle Z(\beta)^{m}\rangle\evaluated{}_{m=0}$. At integer $m>0$, this is an $m$-boundary calculation which can be done using the gravitational path integral and involves wormholes connecting different boundaries. This calculation was attempted [3, 10], but the analytic continuation in $m$ is complicated, and so far, there is no general method to compute $S_{Q}$ in gravitational theories. Recently, [4], among other interesting results, proposed a calculation of $S_{Q}$ using the RMT dual to JT gravity. The interesting proposal was that a new saddle, a one-eigenvalue instanton, dominates in the RMT due to the insertion of the operator $\langle Z(\beta)^{m}\rangle$ in Eq. (2) for large $\beta$ [^2] This new saddle in the matrix integral can potentially be described by a bulk saddle involving a semiclassical brane dual to the one-eigenvalue instanton. Such a description is enticing since the properties of the black hole microstate ground state (in a sector of fixed charge) can be captured through a semiclassical description [4]. The one-eigenvalue instanton indeed correctly computes $\langle Z(\beta)^{m}\rangle$ and, consequently, the annealed entropy. However, we will point out issues with the replica limit when calculating the quenched entropy using this technique.
@@
To ameliorate the situation, we are motivated to study other entropic quantities in this regime. We consider various ensemble-averaged Rényi entropies for the thermal density matrix $\rho=e^{-\beta H}$, summarized in Table 1. As with conventional entropies, the quenched Rényi entropy is positive while the annealed is not sign-definite; as before, evaluating the annealed entropy is simple, whereas the quenched entropy requires taking a delicate replica limit. Due to the subtleties in this replica limit, we propose a new intermediate quantity: the semi-quenched Rényi entropy $S_{SQ}^{(n)}$, which is positive like the quenched Rényi entropy and nevertheless, does not require a replica limit. We explain all these quantities in detail in Sec. II.
@@
In Sec. III, we compute $S_{SQ}^{(n)}$ in JT gravity in the interesting regimes where $S_{A}<0$. This low-temperature regime is generally difficult to understand since the genus expansion in JT gravity breaks down. In Sec. III.1, we sum over all wormhole contriubtions to demonstrate positivity of $S_{SQ}^{(n)}$ in the so-called Airy limit when $\beta=O(e^{2S_{0}/3})$. In Sec. III.2, we show that, when $\beta\gg~O(e^{2S_{0}/3})$, the one-eigenvalue instanton proposed in [4] gives the correct result for $S_{SQ}^{(n)}$. Moreover, we show that these two calculations agree in their overlapping regime of validity. This is our first important result.
@@
Our second goal is to clarify the behavior of the quenched entropy in this low-temperature regime. In Sec. IV we show that while the one-eigenvalue instanton provides the correct answer for $\langle Z(\beta)^{m}\rangle$ for $m\neq 0$, the analytic continuation to $m=0$ required for the quenched entropy is subtle: in the limit $m\to 0$, the one-eigenvalue instanton reunites with the rest of the spectrum and the saddle-point approximation in its derivation breaks down due to a noncommuting order of limits between $S_{0}\to\infty$ and $m\to 0$. We demonstrate that the one-eigenvalue instanton answer disagrees with previous matrix model calculations of the quenched entropy in the low temperature limit [11, 12]. Thus, the one-eigenvalue instanton provides the correct result for computationally simpler quantities like the annealed and semi-quenched entropy, but not for the quenched entropy.
@@ II A zoo of entropies
The different generalizations of the Rényi entropy differ in their ordering of operations: taking logarithms, powers, and ensemble averages. Different choices lead to different quantities, with distinct physical interpretations and computational challenges. Conventionally, the two quantities discussed in the literature are the quenched and annealed entropies. Here, we discuss a larger class of quantities, explaining how each is constructed using a replica trick. These quantities are compared in Table 1.
@@
$=\left(1-n\partial_{n}\right)\log\langle Z_{n}\rangle\evaluated{}_{n=1},$ (7b)
@@
$S_{Q}^{(n)}\equiv\frac{1}{1-n}\left\langle\log\frac{Z_{n}}{Z_{1}^{n}}\right\rangle,$ (8)
@@
$S_{Q}^{(n)}=\frac{1}{1-n}\lim_{\begin{subarray}{c}m\to 0\\ k\to-n\end{subarray}}\left\langle\partial_{m}\left[\left(Z_{n}Z_{1}^{k}\right)^{m}\right]\right\rangle,$ (9)
which can be computed at positive integer $m,\,n,\,k$ using the gravitational path integral and requires analytic continuation for the two replica limits shown above. In the von Neumann limit ($n\to 1$), this gives
@@
$=\left(1-n\partial_{n}\right)\langle\log Z_{n}\rangle\evaluated{}_{n=1},$ (10b)
@@
This quantity can be easily seen to be positive since $Z_{n}\leq Z_{1}^{n}$ for $n\geq 1$ for any member of the ensemble. Taking an ensemble average of this inequality results in the positivity of $S_{SQ}^{(n)}$. For $n<1$, the argument is similar except that the inequalities are reversed.
Finally, in the $n\to 1$ limit, we find a new quantity:
@@
which we call the semi-quenched entropy.  [^3] Another possibility is to define a quasi-quenched Rényi entropy, $S_{QQ}^{(n)}$ $=\frac{1}{1-n}\log\left\langle\frac{Z_{n}}{Z_{1}^{n}}\right\rangle$ $=\frac{1}{1-n}\log\lim_{k\to-n}\left\langle Z_{n}Z_{1}^{k}\right\rangle.$ This still requires a replica limit to compute, and is positive by a similar argument used above. In the $n\to 1$ limit, it yields the quenched entropy. Since it is as difficult to compute as the quenched entropy, we will not focus on it in this paper.
@@ III.1 Airy Wormholes at $\beta=O(e^{2S_{0}/3})$
Figure 2: Comparing $\text{disk}_{\text{res}}^{2}$ and $\text{cyl}_{\text{res}}$, setting $\gamma=1$. At high temperatures (large $\alpha^{-1}$), the disconnected geometry $\text{disk}_{\text{res}}^{2}$ dominates. At low temperatures (small $\alpha^{-1}$), the connected geometry $\text{cyl}_{\text{res}}$ dominates.
We will now use the above idea to compute the semi-quenched Rényi entropy, focusing on the $n=2$ case where we can give explicit formulas. In Appendix A, we show that for all integer $n$, $S_{SQ}^{(n)}\to 0$ as $\alpha\to\infty$. We have
@@
[…]
@@
From Eq. (17), we see that at large $\alpha$, the cylinder contribution dominates over the disk in $\langle Z(\beta)^{2}\rangle$ as shown in Fig. 2 [21]. It is these wormholes that differentiate $S_{SQ}^{(2)}$ from $S_{A}^{(2)}$—which is simply given by the first term in Eq. (18)—rescuing positivity. The fact that $\text{cyl}_{\text{res}}\left(\beta,\beta\right)$ approaches $\text{disk}_{\text{res}}\left(2\beta\right)$ at large $\beta$ is a highly non-trivial consistency condition on wormhole contributions to the gravitational path integral, satisfied by JT gravity.
@@ III.2 Eigenvalue Instantons at $\beta\gg O(e^{2S_{0}/3})$
We discuss aspects of the general solution in Appendix B. To compare the instanton answer with the Airy answer, we focus on the regime $\beta=\alpha e^{2S_{0}/3}$ with $\alpha=O(1)$ but large. Since the remaining eigenvalues remain in their original saddle, we will focus on the effective action for the lowest eigenvalue given by
@@
Eq. (22) implies that $\exp\left(-I[\lambda_{0}^{*}]\right)$ agrees with the leading behaviour of Eq. (17). To compute $S_{SQ}^{(2)}$, we need to find the first subleading correction to the instanton action. In Appendix C, we argue that this comes from a two-eigenvalue instanton that gives us
@@ IV Comments on Quenched Entropy
At $\beta\gtrsim e^{2S_{0}/3}$, the answer for the quenched entropy is also controlled by the Airy limit [11, 12]. Due to the replica limits involved in this calculation, it is much harder to obtain directly but a matrix model analysis was done in [11] where they found that the low temperature entropy behaves like $T^{1+\upbeta}$, where $\upbeta=1,2,4$ is the Dyson index that characterizes the GOE, GUE and GSE ensembles  [^7] This behavior holds at arbitrarily low temperatures, up to doubly non-perturbative corrections [11].. In particular, for GUE, the leading order term at large $\alpha$ is $S_{Q}^{(n)}\approx\frac{7\pi^{4}\gamma^{3}}{360\alpha^{3}}\frac{(1+n)(1+n^{2})}{n^{3}}$. The quenched entropy, obtained numerically by sampling 1000 GUE $\times 1000$ 1000 matrices, is plotted in Fig. 1 and matches the aforementioned analytic result at large $\alpha$.
In contrast, the proposal of using the one-eigenvalue instanton for the quenched entropy [4] does not reproduce this, leading to $S_{Q}\approx T^{3/2}$ at low temperatures, independent of the symmetry class.
As seen above, the location of the instanton is a function of $m$ and, unlike the semi-quenched entropy, approaches the continuum edge as we take $m\to 0$. In this limit, the saddle point approximation breaks down since higher-loop contributions in Eq. (22) become important. One should have always been suspicious of this analysis since the leading answer comes from the continuum spectrum in this limit, whereas the assumption used in obtaining the instanton was that the leading eigenvalue dominates $Z\left(\beta\right)$. This is a consequence of taking the wrong order of limits between $\beta\to\infty$ and $m\to 0$  [^8] This is similar in spirit to issues arising for Rényi entropies for states where fluctuations in the geometry are large [22, 23]..

### v4

@@ Abstract
In Jackiw-Teitelboim (JT) gravity, which is dual to a random matrix ensemble, the annealed entropy differs from the quenched entropy at low temperatures and goes negative. However, computing the quenched entropy in JT gravity requires a replica limit that is poorly understood. To circumvent this, we define an intermediate quantity called the semi-quenched entropy, which has the positivity properties of the quenched entropy, while requiring a much simpler replica trick. We compute this in JT gravity in different regimes using i) a bulk calculation involving wormholes corresponding to the Airy limit of the dual matrix integral and ii) a boundary calculation involving one-eigenvalue instanton saddles proposed by Hernández-Cuenca, demonstrating consistency between these two calculations in their common regime of validity. We also clarify why similar one-eigenvalue instanton saddles cannot be used to compute the quenched entropy due to a breakdown of the saddle-point approximation for the one-eigenvalue instanton in the replica limit. Our results show how to use the gravitational path integral to prove that black holes in JT gravity have isolated ground states and to study their properties.
@@ I Introduction
Do black holes have isolated ground states? In this paper, we shall answer this question using a gravitational calculation in a two-dimensional toy model, JT gravity, which has served in recent years as a tool for understanding quantum gravity effects in realistic black holes. As a two-dimensional toy model of quantum gravity, it is computationally tractable while still exhibiting interesting non-perturbative effects. The action of this theory is given by [1]
@@
In [2], the Euclidean gravitational path integral of JT gravity was computed exactly to all perturbative orders in $e^{-S_{0}}$, in the $S_{0}\to\infty$ limit. The partition function at finite temperature $Z(\beta)$ is computed by fixing the length of the boundary and summing over all hyperbolic geometries with different topologies satisfying the boundary conditions. The resulting asymptotic genus expansion agrees with the double-scaling limit of a random matrix theory (RMT):
@@
where $H$ is a Hermitian matrix and the potential $V(H)$ is specified by the leading-order density of states:
@@
Figure 1: Comparing semi-quenched, annealed and quenched second Rényi entropies, zoomed in near the Airy edge $\beta=\alpha e^{2S_{0}/3}$ (with $\gamma=1$). Semi-quenched and annealed Rényis are obtained analytically, the quenched Rényi is obtained numerically. Notice the different low temperature behavior (inset): the quenched entropy has a $\alpha^{-3}$ power-law decay which can be obtained analytically (dotted blue line), whereas the semi-quenched entropy decays exponentially. Unlike the annealed entropy, both these entropies remain positive.
@@
In the absence of matter fields, the only observable in pure JT gravity is the energy spectrum. The obvious quantity of interest that probes our initial question is the thermal entropy, obtained from the thermal partition function $Z(\beta)$. If the thermal entropy vanishes at small temperatures (for which we probe the ground state), then black holes would indeed have a single isolated ground state, separated from the rest of the spectrum by a gap. However, in an ensemble of theories where one only computes averaged quantities, one has to be careful about which entropy is being computed, since the average can be taken in different ways. Conventionally, the quantities discussed in the literature are the annealed entropy and the quenched entropy. The annealed entropy $S_{A}\left(\beta\right)\equiv\left(1-\beta\partial_{\beta}\right)\log\langle Z\left(\beta\right)\rangle$ differs from the quenched entropy, $S_{Q}\left(\beta\right)\equiv\left(1-\beta\partial_{\beta}\right)\langle\log Z\left(\beta\right)\rangle$, which is truly the average entropy of the ensemble. An important distinction between these entropies is that the quenched entropy is positive, while the annealed entropy can be negative.
@@
In gravitational theories, the annealed entropy is much easier to compute. When the entropy is self-averaging, $S_{A}=S_{Q}$. However, in JT gravity, $S_{A}$ fails to be self-averaging at sufficiently low temperatures. The consequent mismatch between the two entropies was analyzed in [3], where they found $S_{A}<0$ at $\beta=O(e^{2S_{0}/3})$ [5]. The negativity of the annealed entropy is not unique to the JT toy model; rather, it persists in higher-dimensional theories for all near-extremal black holes that are not supersymmetric at extremality [6, 7, 8, 9]. This low temperature regime is thus among the most promising to probe the physics of only a small number of black hole microstates near the ground state. To address our initial question, we need to instead understand how the quenched entropy behaves in this regime.
[…]
@@
To ameliorate the situation, we are motivated to study other entropic quantities in this regime. We consider various ensemble-averaged Rényi entropies for the thermal density matrix $\rho=e^{-\beta H}$, summarized in Table 1. As with conventional entropies, the quenched Rényi entropy is positive while the annealed is not sign-definite; as before, evaluating the annealed entropy is simple, whereas the quenched entropy requires taking a delicate replica limit. Due to the subtleties in this replica limit, we propose a new intermediate quantity: the semi-quenched Rényi entropy $S_{SQ}^{(n)}$, which is positive like the quenched Rényi entropy and, nevertheless, does not require a replica limit. While the semi-quenched entropy might be easy to compute, it is as powerful as the quenched entropy in answering our starting question. In fact, we show that if the semi-quenched entropy remains positive for all temperatures, then the ground state of the spectrum is indeed isolated and the quenched entropy, even if not computable analytically, is also provably positive.
@@
In Sec. III, we compute $S_{SQ}^{(n)}$ in JT gravity in the interesting regimes where $S_{A}<0$. This low-temperature regime is generally difficult to understand since the genus expansion in JT gravity breaks down. In Sec. III.1, we sum over all wormhole contributions to demonstrate positivity of $S_{SQ}^{(n)}$ in the so-called Airy limit when $\beta=O(e^{2S_{0}/3})$. In Sec. III.2, we show that, when $\beta\gg~O(e^{2S_{0}/3})$, the matrix integral dual to JT gravity is dominated by a new saddle, the one-eigenvalue instanton proposed in [4], in which the lowest eigenvalue is well-separated from the rest of the eigenvalues. This saddle gives the correct positive result for $S_{SQ}^{(n)}$. Moreover, we show that these two calculations agree in their overlapping regime of validity. The positivity of the semi-quenched entropy for all values of $\beta$ leads us to the conclusion that we have an isolated ground state in each member of the ensemble. This is our first important result.
@@
Our second goal is, for completeness, to clarify the behavior of the quenched entropy in the low-temperature regime. In Sec. IV we show that while the one-eigenvalue instanton provides the correct answer for $\langle Z(\beta)^{m}\rangle$ for $m\neq 0$, the analytic continuation to $m=0$ required for the quenched entropy is subtle: in the limit $m\to 0$, the one-eigenvalue instanton reunites with the rest of the spectrum and the saddle-point approximation in its derivation breaks down due to a noncommuting order of limits between $S_{0}\to\infty$ and $m\to 0$. We demonstrate that the one-eigenvalue instanton answer disagrees with previous matrix model calculations of the quenched entropy in the low temperature limit [11, 12]. Thus, the one-eigenvalue instanton provides the correct result for computationally simpler quantities like the annealed and semi-quenched entropy, but not for the quenched entropy.
@@ II A zoo of entropies
The different generalizations of the Rényi entropy differ in their ordering of operations: taking logarithms, powers, and ensemble averages. Different choices lead to different quantities, with distinct physical interpretations and computational challenges. Conventionally, the two quantities discussed in the literature are the quenched and annealed entropies. Here, we discuss a larger class of quantities, summarized in Table 1.
@@
[…]
@@
$S_{Q}^{(n)}\equiv\frac{1}{1-n}\left\langle\log\frac{Z_{n}}{Z_{1}^{n}}\right\rangle=\frac{\left\langle\log Z_{n}\right\rangle-\left\langle\log Z_{1}^{n}\right\rangle}{1-n},$ (8)
@@
$S_{Q}^{(n)}=\frac{1}{1-n}\lim_{m\to 0}\left[\frac{\left\langle\left(Z_{n}\right)^{m}\right\rangle-\left\langle\left(Z_{1}\right)^{nm}\right\rangle}{m}\right],$ (9)
which can be computed at positive integer m,n,km,\,n,\,k using the gravitational path integral and requires analytic continuation for the two replica limits shown above. In the von Neumann limit ($n\to 1$), this gives
@@
[…]
@@
In the $n\to 1$ limit, we find a new quantity:
@@
which we call the semi-quenched entropy.  [^3] Another possibility is to define a quasi-quenched Rényi entropy, $S_{QQ}^{(n)}$ $=\frac{1}{1-n}\log\left\langle\frac{Z_{n}}{Z_{1}^{n}}\right\rangle$ $=\frac{1}{1-n}\log\lim_{k\to-n}\left\langle Z_{n}Z_{1}^{k}\right\rangle.$ This still requires a replica limit to compute, and is positive by a similar argument used above. In the $n\to 1$ limit, it yields the quenched entropy. Since it is as difficult to compute as the quenched entropy (which also requires one replica limit), we will not focus on it in this paper.
These entropies can be easily seen to be positive since $Z_{n}\leq Z_{1}^{n}$ for $n\geq 1$ for any member of an ensemble of theories with discrete eigenenergies. Taking an ensemble average of this inequality results in the positivity of $S_{SQ}^{(n)}$. For $n<1$, the argument is similar except that the inequalities are reversed. An important fact is that $Z_{n}>Z_{1}^{n}$ for $n\geq 1$ as $T\to 0$ for any member of the ensemble for which the ground state is not isolated from the rest of the spectrum (see Appendix A); for all other members $Z_{n}-Z_{1}^{n}$ decreases with $T$ and vanishes as $T\to 0$. Therefore, if there is even one member with a non-isolated ground state, this would be reflected in the semi-quenched entropy eventually becoming negative. Consequently, positivity of the semi-quenched entropy implies an isolated ground state for all members of the ensemble. Positivity of the quenched entropy also follows, due to its positivity in each ensemble member.
@@ III.1 Airy Wormholes at $\beta=O(e^{2S_{0}/3})$
We will now use the above idea to compute the semi-quenched Rényi entropy, focusing on the $n=2$ case where we can give explicit formulas. In Appendix B, we show that for all integer $n$, $S_{SQ}^{(n)}\to 0$ as $\alpha\to\infty$. We have
@@
where $\text{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}dt\,e^{-t^{2}}$. Using these formulas, we obtain:
@@
From Eq. (17), we see that at large $\alpha$, the cylinder contribution dominates over the disk in $\langle Z(\beta)^{2}\rangle$ [21]. It is these wormholes that differentiate $S_{SQ}^{(2)}$ from $S_{A}^{(2)}$—which is simply given by the first term in Eq. (18)—rescuing positivity. The fact that $\text{cyl}_{\text{res}}\left(\beta,\beta\right)$ approaches $\text{disk}_{\text{res}}\left(2\beta\right)$ at large $\beta$ is a highly non-trivial consistency condition on wormhole contributions to the gravitational path integral, satisfied by JT gravity.
@@ III.2 Eigenvalue Instantons at $\beta\gg O(e^{2S_{0}/3})$
We discuss aspects of the general solution in Appendix C. To compare the instanton answer with the Airy answer, we focus on the regime $\beta=\alpha e^{2S_{0}/3}$ with $\alpha=O(1)$ but large. Since the remaining eigenvalues remain in their original saddle, we will focus on the effective action for the lowest eigenvalue given by
@@
Eq. (22) implies that $\exp\left(-I[\lambda_{0}^{*}]\right)$ agrees with the leading behaviour of Eq. (17). To compute $S_{SQ}^{(2)}$, we need to find the first subleading correction to the instanton action. In Appendix D, we argue that this comes from a two-eigenvalue instanton that gives us
@@
We have thus proven positivity of the semi-quenched entropy for any value of $\beta$, and therefore that all members of the ensemble have an isolated ground state. In turn, this implies that the gravitational quenched entropy also stays positive for any value of $\beta$.
@@ IV Comments on Quenched Entropy
At $\beta\gtrsim e^{2S_{0}/3}$, the answer for the quenched entropy is also controlled by the Airy limit [11, 12]. Due to the replica limits involved in this calculation, it is much harder to obtain directly but a matrix model analysis was done in [11] where they found that the low temperature entropy at leading order term at large $\alpha$ is $S_{Q}^{(n)}\approx\frac{7\pi^{4}\gamma^{3}}{360\alpha^{3}}\frac{(1+n)(1+n^{2})}{n^{3}}$. [^7] This behavior holds at arbitrarily low temperatures, up to doubly non-perturbative corrections [11]. The quenched entropy, obtained numerically by sampling 1000 GUE $1000\times 1000$ matrices, is plotted in Fig. 1, matching the aforementioned analytic result at large $\alpha$.
In contrast, the proposal of using the one-eigenvalue instanton for the quenched entropy [4] does not reproduce this, leading to $S_{Q}\approx T^{3/2}$ at low temperatures, independent of the symmetry class. The location of the instanton is a function of $m$ and approaches the continuum edge as we take $m\to 0$. In this limit, the saddle point approximation breaks down since higher-loop contributions in Eq. (22) become important. This is also problematic because the leading answer for quenched in [4] comes from the continuum spectrum, whereas the assumption used in obtaining the instanton was that the leading eigenvalue dominates $Z\left(\beta\right)$. This is a consequence of a wrong order of limits between $\beta\to\infty$ and $m\to 0$  [^8] This is similar in spirit to issues arising for Rényi entropies for states where fluctuations in the geometry are large [22, 23]..